OpenStudy (anonymous):

hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won championship with 51 points. they won 9 more games than they tied. How many win and ties did the team have?

OpenStudy (anonymous):

tied 11 games. won 20 2(x+9) +1(x) = 51 2 for the points for a win, x+9 because they won 9 more games than they tied, x = games tied, and 1 for the points for a tie. 2x+18+x=51 3x=33 x=11=games tied, 11+9 = games won